### v2

@@ IX Radiometeorology Mixed URSI, IUGG, IUPAP
The second line of action was closely linked to the strategy of internationalization of the society’s journals. From the outset, it was planned that the papers from the conference would be featured in a UNESCO-sponsored special issue, to be published in Il Nuovo Cimento’s supplement, known as Supplemento al Nuovo Cimento. [^56] This was the issue number 2 of Supplemento al Nuovo Cimento of Vol. 6 of Il Nuovo Cimento published in 1949, see https://link.springer.com/journal/40761/volumes-and-issues/6-2/supplement. The SIF Council saw the special issue as a way to strengthen and quicken the growth and international standing of Il Nuovo Cimento. [^57] See, Minutes of the SIF General Assembly, November 9, 1948, ASIF. Polvani, along with other SIF Council members, championed this internalization effort despite facing opposition from some of the society’s more traditionalist members, who lamented the excessive use of English in the flagship journal of the Italian Physical Society. [^58] Minutes of the SIF General Assembly, September 16, 1950; and Minutes of the SIF Council Meeting, October 15, 1950, ASIF.
@@
The third and perhaps more subtle initiative involved the concurrent planning of a separate event: an international conference on cosmic rays, scheduled for September 1949 in Como, Alessandro Volta’s birthplace, to celebrate the 150th anniversary of the invention of the Voltaic pile. This conference was sponsored by the second IUPAP topical commission established in 1947, the Commission on Cosmic Rays. From its inception, the conference was envisioned to complement an IUPAP-sponsored international congress in Basel, so that the joint venture could attract to Como many “world-famous non-Italian scientists." [^59] Minutes of the SIF General Assembly, November 9, 1948, ASIF. Especially notable was the anticipated attendance of Enrico Fermi, who had never returned to Europe after leaving Italy in 1938 [3]. Given Fermi’s planned visit to Italy and Europe, the SIF Council regarded the Como conference as potentially surpassing the Florence conference in significance and scope. By the end of 1948, SIF began collaborating with IUPAP on both the Florence and Como conferences, enhancing their mutual impact and showcasing the Italian physics community as exceptionally active in hosting IUPAP-affiliated international scientific gatherings. The strategic linkage and synergy between these two conferences, as revealed in Polvani’s uncatalogued correspondence stored in the SIF archives, underscore their integral role in the broader strategy to internationalize Italian physics, with IUPAP serving as a key global partner during this period. [^60] See, numerous documents dated 1948/49 in Polvani-ASIF.
@@
These activities underscore the complex interplay of strategies deployed by leading Italian physicists to elevate the stature of Italian physics through concerted internationalization efforts. The Florence conference emerged as a pivotal moment in this ambitious campaign, marking the first significant step by SIF and the Italian physics community to reassert and reclaim a prominent position within the international scientific arena in the aftermath of World War II. This strategy proved so effective that by 1950, Amaldi could proudly declare that “Italy was the favored nation" for hosting IUPAP-endorsed international conferences and congresses in the immediate postwar era, a testament to the success of Italian physicists’ approach. [^61] Minutes of the SIF Council Meeting, September 16, 1950, ASIF.
@@ 6 The role of the Florence conference in an emerging research field
The conference took place over four days, from May 17 to 20, 1949, with the opening ceremony hosted at Villa Favard, at that time the auditorium for the Faculty of Economic and Commercial Sciences of the University of Florence, and the scientific sessions predominantly conducted at the National Museum of the History of Science, now known as Museo Galileo (see program in Fig. 4). [^62] Borghi, Press release, International Conference on Statistical Mechanics, May 16, 1949, 19/c-ASUFi. The last scientific session and the closing gala dinner occurred at the Albergo Ristorante Aurora in Fiesole. Official records indicate a participation of seventy to eighty attendees, with a notable attendance sheet featuring signatures of distinguished participants (Fig. 5).
@@
To capture how the Commission was reconfiguring its focus and scope through the Florence conference it is instructive to make a comparison with the 1948 symposium held in Brussels. The Florence conference notably surpassed the Brussels symposium in international participation, attracting nearly triple the number of attendees and featuring speakers from eight diverse countries (see Appendix A), as opposed to the Brussels symposium, where speakers hailed predominantly from five Western European nations, with the majority originating from just Belgium (7) and The Netherlands (4) (see list of participants in [48]). Despite the number of speakers being very similar in the two conferences (around fifteen), the Florence event demonstrated a significantly broader geographical distribution. This expansion in reach, while primarily within the Euro-Atlantic sphere akin to the IUPAP membership of that era, nonetheless marked a considerable enlargement of the Commission’s international engagement. Notably, the Florence conference included participants from two continents and featured at least one delegate from the emerging Eastern bloc, the Polish physicist Jan Weyssenhoff, underscoring the conference’s role in broadening the Commission’s international footprint.
@@
Scientifically, the distinct emphases of the two meetings are clearly underscored by their titles. The Brussels symposium was dedicated to thermodynamics, in contrast to the Florence conference, which centered on Statistical Mechanics. This distinction goes much beyond differences between titles. It was significantly reflected in the themes of the presentations, as documented in their respective Proceedings. The presentations at Brussels were signaling that the main focus of research concerning the Commission was statistical thermodynamics. The animated discussions following most presentations at the Brussels symposium showed that many of the employed assumptions and approaches were matter of contention. This was particularly evident in the opening scientific talk,“The statistical basis of thermodynamics," which explored the foundation of thermodynamic laws on a unique statistical principle [22]. [^66] Guggenheim was also present at the Florence conference, where he gave a talk discussing the Boltzmann factor [21]. Given by the President of the Commission, Edward Guggenheim, the opening talk of the symposium suggested a collective lean towards recognizing statistical mechanics as a vital area for developing an international research community. While, for obvious reasons, in the Brussels’ proceedings the lion’s share is taken up by the thermodynamics of open systems, in Florence several contributions focused on more fundamental problems of statistical mechanics [12, 29, 21, 10, 27]; this passage accompanied a more basic Commission’s shift in the scientific direction, prefigured by its renaming. As discussed in the Section 4, moving from Commission on Thermodynamic Magnitudes and Notations to Commission on Thermodynamics and Statistical Mechanics had already marked the progression toward assuming a scientific role rather than a purely technical one as previously envisaged. This strategic shift to statistical mechanics, seen as closer to the heart of physics and offering a richer vein of research possibilities, was evident in the content featured at the Florence conference in contrast to the Brussels symposium.
@@
The Commission’s shift toward statistical mechanics was operationalized by the Florence conference through the delineation of the five above-mentioned sub-topics. This strategic focus underscored the Commission’s evolution beyond its traditional thermodynamics realm, aiming instead to anchor itself within the physics community by spearheading the exploration of specific, cutting-edge research areas. The commitment to these sub-topics signified an intention not just to serve as a community hub but to actively shape the research landscape. This ambition was realized as the proceedings of the conference showcased: all five sub-topics were comprehensively addressed by the speakers, albeit with a variable number of papers dedicated to each sub-topic. [^67] For the mathematical methods of the configuration integral, see [39]; cooperative phenomena were addressed in [21, 51]; distribution functions were the topic of [37, 49]; the statistical mechanics of irreversible phenomena was explicitly tackled by [13, 27, 59]; and the quantum foundations of statistical mechanics was the topic of Born’s lecture [12]. Such thorough coverage reaffirms the conference’s success in redirecting the attention of the community around the Commission towards statistical mechanics. The opening speech of Kramers [30] openly reveals the intention to position statistical mechanics as a field with historical roots in the kinetic theory of gases, now poised to incorporate the recent advances in quantum physics and to confront fundamental and outstanding problems, including the derivation of the thermodynamic properties of a system from the interaction among its elementary constituents, starting from the comprehension of ferromagnetism and the condensation of a gas.
@@
Some presentations at the conference have been recognized, a posteriori, as influential on the later developments of what was then considered an emerging field. This is particularly manifest in Lars Onsager’s lecture, titled Statistical Hydrodynamics. His paper published in Supplemento al Nuovo Cimento [44] has amassed over 2,000 citations to date, dwarfing the citation counts of other papers from the proceedings by two orders of magnitude. [^68] This analysis was made using Google Scholar. A research on Scopus confirms the later impact of the paper with a count of 1362 citations from 1970 to 2024 with an evident increase of citations starting from the late 1990s. Onsager’s paper studies the statistics of a system of vortices, that represents the first and one of the most important examples displaying (absolute) negative temperatures. Remarkably, at the time, the significance of this work was not fully recognized, just as it is notable that the discovery for which Onsager was awarded the Nobel Prize for Chemistry in 1968 (his reciprocal relations identified in 1931 [42, 43]) was the focus of Casimir’s presentation at the Florence Conference, rather than his own [13].
@@ 7 The solidification of a tradition from topical conferences to StatPhys
1 Florence (Italy) [^1] Convegno internazionale di meccanica statistica 1949 16 Boston (Usa) 1986
@@ 15 Edinburgh (UK) 1983
7070footnotetext: This table was compiled using data sourced from the IUPAP reports, Series B2aa “General Reports,” Vols.1-3, IuG. The denomination StatPhys was officially introduced in 1977 to designate the Haifa conference as StatPhys13. By assigning this sequence number, the Commission implicitly applied the StatPhys label to all prior meetings in the series, thereby retrospectively recognizing them as part of the StatPhys conference lineage, as shown in the table.
@@
The 1970s witnessed significant transformations concerning the policies and structures of IUPAP commissions and their sponsored conferences. Starting from the 1969 General Assembly in Dubrovnik, IUPAP embarked on a comprehensive revision of policies related to the composition of commissions and the criteria for conference sponsorship. In response to the increasing volume of conference proposals from IUPAP commissions, the Executive Committee of IUPAP introduced a set of evaluation criteria focused on scientific merit, international representation, and organizational solidity. A member of the U.S. national committee, William Havens Jr., further contributed to these deliberations by proposing a taxonomy for classifying conferences into three distinct categories, which received tentative approval during the 13th General Assembly in Dubrovnik. The type-A conferences, called “general conferences," were designed “to provide an overview of the entire field of interest to a Commission, and would normally occur at three-year intervals" with a foreseeable attendance of about 750-1500 scientists. Type-B conferences were called “topical conferences" and were designed to “concentrate on broad sub-fields" within the general of the Commission’s interest. The foreseen attendance was about 300-600 physicists. The smaller Type C conference were called “special conferences" and “would concentrate on much more restricted specialized topics" than the type-B conferences. The expected attendance was in the range of 50-200 physicists. [^71] IUPAP, Report 1970, esp. p. 25, p. 29, pp. 81-83, Series B2aa, Vol. 2, IuG. At the subsequent IUPAP General Assembly, held in Washington D.C. in 1972, new procedural guidelines were established for the composition of commissions and the election of their members. Incidentally, a new identification system for commissions was adopted, denoted as ‘C. followed by a sequential number reflecting their chronological establishment, [^72] IUPAP, Report 1973, pp. 14-18, Series B2aa, Vol. 2, IuG. a system that remains in use today.’ [^73] The full point after C was later eliminated and currently IUPAP commissions are simply denoted by C plus a sequential number, with C3 being the denomination of the Commission on Statistical Physics.
@@
The reorganization within IUPAP had significant implications for the operations of all its commissions. In 1972, the designation of the Commission on Thermodynamics and Statistical Mechanics as C.3 cemented its reputation as the inaugural topical commission of IUPAP, following C.1, the Commission on Finances, and C.2, the general-purpose SUN Commission. Prompted by these general reconfigurations within IUPAP, in 1973 the Commission resolved to inaugurate an award recognizing outstanding achievements in thermodynamics and statistical mechanics. The newly appointed chairman, Canadian physicist Donald D. Betts, announced to the Commission members that the concept of the award was met with enthusiastic endorsement, both within the commission and among the wider physics community. Betts advocated for the award to take the form of a gold medal rather than a cash prize, arguing that it would not only be economically feasible for the Commission but also offer a lasting symbol of achievement. Following a Commission vote on names of founding fathers of the sub-discipline, the award was called the Boltzmann Medal, in honor of one of the foundational figures in the discipline. The inaugural Boltzmann Medal was presented at the 1975 conference on Statistical Mechanics to Kenneth Wilson, recognizing his pioneering work on the renormalization group. This decision came despite Betts’ earlier communication in 1973, which had suggested Lars Onsager as the “obvious choice" [^74] Betts to Members of the Commission on Thermodynamics and Statistical Mechanics, March 9, 1973, see also Betts to Larkin Kerwin, September 19, 1973, Series E3 “Correspondence of the Commissions,” Vol. 2, Folder C 3, IuG. for his contributions to the study of irreversible processes, for which he had already been awarded the Nobel Prize in Chemistry in 1968.757575Lars Onsager died in 1976, it was therefore not possible to take him into consideration as recipient of the next Boltzmann Medals. Later, in 1982, Kenneth Wilson also won the Nobel Prize in Physics. The same occurred to Giorgio Parisi who received the Boltzmann medal in 1992 and the Nobel prize in 2021.
@@
The second significant shift in the mid-1970s aimed to establish a more systematic scheduling of the Commission on Thermodynamics and Statistical Mechanics’ international conferences, specifically, to formalize a biennial interval for these conferences starting in 1975.767676Betts to Kerwin, September 19, 1973, Series E3 “Correspondence of the Commissions,” Vol. 2, Folder C 3, IuG. This initiative was in line with above-mentioned broader initiative to categorize IUPAP-sponsored international conferences, as endorsed during the 1969 IUPAP General Assembly. The conferences organized by Commission C.3 naturally fell into the Type-A category, embodying the essence of a general conference. The decision in the early 1960s to forgo specific titles in favor of a more inclusive approach was exactly in this direction. However, aligning the Commission’s ambition for biennial conferences with IUPAP’s organizational framework proved challenging. The financial and administrative constraints of adhering to a two-year cycle were compounded by IUPAP’s triennial budgetary cycle, which spanned from one General Assembly to the next. As Larkwin Kerwin, the IUPAP General Secretary, explained to Betts, supporting Type-A conferences biennially within the confines of a three-year budget was problematic. [^77] Kerwin to Betts, October 22, 1973, Series E3, Vol. 2, Folder C 3, IuG. Additionally, during this period, IUPAP officials urged all commissions to reassess their roles in light of recent disciplinary advancements. Commission C.3 was tasked to provide a report in 1977 “examin[ing] the boundaries of its field and to consider whether these should be modified, whether the Commission should be divided, fused with another commission, or dissolved" as well as to reassess its field through the lens of its international conferences. [^78] Kerwin to Herbert B. Callen, October 10, 1975, Series E3, Vol. 2, Folder C 3, IuG. This directive underscored the need for commissions to remain adaptive and responsive to the evolving landscape of their respective disciplines.
@@
This request initiated a self-reflection among the members of Commission C.3 regarding its activities, roles, and functions. In 1977, the chair of the Commission in 1975-78, U.S. statistical physicist Herbert B. Callen, submitted a preliminary report to Kerwin, articulating the Commission members’ perception of both the Commission and the field it served during that time. It is beneficial to quote extensive passages from this report, as they directly illustrate the views of Commission members about the Commission itself and its scientific domain within the broader landscape of physics.

### v3

@@ IX Radiometeorology Mixed URSI, IUGG, IUPAP
The second line of action was closely linked to the strategy of internationalization of the society’s journals. From the outset, it was planned that the papers from the conference would be featured in a UNESCO-sponsored special issue, to be published in Il Nuovo Cimento’s supplement, known as Supplemento al Nuovo Cimento. [^63] This was the issue number 2 of Supplemento al Nuovo Cimento of Vol. 6 of Il Nuovo Cimento published in 1949, see https://link.springer.com/journal/40761/volumes-and-issues/6-2/supplement. The SIF Council saw the special issue as a way to strengthen and quicken the growth and international standing of Il Nuovo Cimento. [^64] See, “Minutes of the SIF General Assembly, November 9, 1948,” ASIF. Polvani, along with other SIF Council members, championed this internalization effort despite facing opposition from some of the society’s more traditionalist members, who lamented the excessive use of English in the flagship journal of the Italian Physical Society.656565“Minutes of the SIF General Assembly, September 16, 1950;” and “Minutes of the SIF Council Meeting, October 15, 1950,” ASIF.
@@
The third and perhaps more subtle initiative involved the concurrent planning of a separate event: an international conference on cosmic rays, scheduled for September 1949 in Como, Alessandro Volta’s birthplace, to celebrate the 150th anniversary of the invention of the Voltaic pile. This conference was sponsored by the second IUPAP topical commission established in 1947, the Commission on Cosmic Rays. From its inception, the conference was envisioned to complement an IUPAP-sponsored international congress in Basel, so that the joint venture could attract to Como many “world-famous non-Italian scientists."666666“Minutes of the SIF General Assembly, November 9, 1948,” ASIF. Especially notable was the anticipated attendance of Enrico Fermi, who had never returned to Europe after leaving Italy in 1938 [46]. Given Fermi’s planned visit to Italy and Europe, the SIF Council regarded the Como conference as potentially surpassing the Florence conference in significance and scope. [^67] During his 1949 trip, Fermi gave a number of lectures on atomic physics in Milan and Rome. For the English translation of of these lectures with commentaries, see [78]. By the end of 1948, SIF began collaborating with IUPAP on both the Florence and Como conferences, enhancing their mutual impact and showcasing the Italian physics community as exceptionally active in hosting IUPAP-affiliated international scientific gatherings. The strategic linkage and synergy between these two conferences, as revealed in Polvani’s uncatalogued correspondence stored in the SIF archives, indicate their integral role in the broader strategy to internationalize Italian physics, with IUPAP serving as a key partner during this period. [^68] See, numerous documents dated 1948/49 in Polvani-ASIF.
@@
These activities underscore the complex interplay of strategies deployed by leading Italian physicists to elevate the stature of Italian physics through concerted internationalization efforts. The Florence conference emerged as a pivotal moment in this ambitious campaign, marking the first significant step by SIF and the Italian physics community to reassert and reclaim a prominent position within the international scientific arena in the aftermath of World War II. This strategy proved so effective that by 1950, Amaldi could proudly declare that “Italy was the favored nation" for hosting IUPAP-endorsed international conferences and congresses in the immediate postwar era, a testament to the success of Italian physicists’ approach.696969“Minutes of the SIF Council Meeting, September 16, 1950,” ASIF.
@@ 6 The role of the Florence conference in an emerging research field
The conference took place over four days, from May 17 to 20, 1949, with the opening ceremony hosted at Villa Favard, at that time the auditorium for the Faculty of Economic and Commercial Sciences of the University of Florence, and the scientific sessions predominantly conducted at the National Museum of the History of Science (see program in Fig. 4). [^70] Borghi, Press release, International Conference on Statistical Mechanics, May 16, 1949, 19/c-ASUFi. The last scientific session and the closing gala dinner occurred at the Albergo Ristorante Aurora in Fiesole. Official records indicate a participation of seventy to eighty attendees, with a notable attendance sheet featuring signatures of distinguished participants (Fig. 5).
@@
In order to contextualize the conference’s role in sub-discipline formation from a scientific standpoint, let us briefly discuss the status of statistical mechanics at the time, noting that historical reconstructions of its development throughout the 20th century are still incomplete. [^74] To our knowledge, the most comprehensive accounts of the long-term development of statistical mechanics are the contributions by Cyril Domb (1995) and Max Dresden (1995) in [14] At the start of the century several scientific sub-fields were interconnected not only because some scientists contributed to multiple domains, but also because it was possible to borrow tools and concepts from other fields. [^75] From this point of view, statistical mechanics has always been strongly interdisciplinary. A notable example is the statistical considerations used to study black body radiation, which led Max Planck to formulate his quantum hypothesis [2].
The 1930s marked a transitional period in both education and research. Statistical ensembles were known from the era’s onset, but Fowler’s 1936 book on statistical mechanics [29] mentions the canonical ensemble only in passing, while a few years later, Schrödinger significantly contributed to the dissemination of Gibbs’ ideas [21]. This period also saw significant advances in the study of phase transitions, perhaps the most important and succesful domain in statistical physics. The existence of transitions between different equilibrium phases of a substance with varying temperature or pressure was well-known, as was the law of corresponding states, an early example of universality [51]. However, precise experiments and Lars Onsager’s 1944 exact solution of the two dimensional Ising model [61] were needed to convince the scientific community that the characterization of a phase transition as described by the van der Waals equation was incorrect.
By the time of the Florence conference, the understanding of equilibrium phenomena was advanced enough to guide research and clarify several open problems [81]: finding techniques to determine the partition function, developing a satisfactory theory of liquids, and moving beyond the classical theory of phase transitions, to name a few. In contrast, nonequilibrium phenomena were much less understood [22]. In addition, they covered such a wide range of phenomena that it was (and still is) impossible to encompass them within a single theory. After the formulation of the Boltzmann equation, a significant advancement was made by Onsager, who described transport processes in systems kept not too far from equilibrium, leading to the development of “nonequilibrium thermodynamics" [59].
Systems far from equilibrium can exhibit a completely different phenomenology. In this domain, Prigogine and his Belgian school were paying a pivotal role. They presented their approach at the 1948 Brussels symposium and in the 1949 Florence. [^76] Prigogine’s work in this period laid the groundwork for his 1960s research on the role of dissipative structures in thermodynamic systems far from equilibrium, which would earn him the Nobel Prize in Chemistry in 1977; for a review of Prigogine’s theory, see [69]. However, Prigogine encountered considerable skepticism when he presented his work on irreversible phenomena at the Brussels symposium. He recalled that his presentation was met with hostility, for the prevailing view among experts was that statistical mechanics should focus on equilibrium states, viewing irreversible processes as merely transitory phenomena [68]. [^77] If we trust Prigogine’s recollection, we might assume that, by designating irreversible phenomena as one of the five central topics of the Florence conference, Prigogine, with the support of Kramers, aimed to challenge this prevailing view and position the study of irreversible phenomena as as cutting-edge research agenda within the community being fostered through such gatherings.
More generally, after World War II, the field of statistical mechanics was not only fragmented into various research trajectories but also deeply intertwined with the evolution of emerging research fields, whose development varied significantly across different countries. In particular, it was connected to the growing research sub-fields of chemical physics and solid-state physics. Regarding Italy, more specifically, statistical mechanics was still in its infancy, with little to no established research agendas actively advancing the field [17].
@@
To capture how the Commission was reconfiguring its focus and scope through the Florence conference it is instructive to make a comparison with the 1948 symposium held in Brussels. The Florence conference notably surpassed the Brussels symposium in international participation, attracting nearly triple the number of attendees and featuring speakers from eight diverse countries (see Appendix A), as opposed to the Brussels symposium, where speakers hailed predominantly from five Western European nations, with the majority originating from just Belgium (7) and The Netherlands (4) [see list of participants in 66]. Despite the number of speakers being very similar in the two conferences (around fifteen), the Florence event demonstrated a significantly broader geographical distribution. This expansion in reach, while primarily within the Euro-Atlantic sphere akin to the IUPAP membership of that era, nonetheless marked a considerable enlargement of the Commission’s international engagement. Notably, the Florence conference included participants from two continents and featured at least one delegate from the emerging Eastern bloc, the Polish physicist Jan Weyssenhoff, underlaying the conference’s role in broadening the Commission’s international footprint.
@@
Scientifically, the distinct emphases of the two meetings are clearly shown by their titles. The Brussels symposium was dedicated to thermodynamics, in contrast to the Florence conference, which centered on statistical mechanics. This distinction goes much beyond differences between titles. It was significantly reflected in the themes of the presentations, as documented in their respective proceedings. The presentations at Brussels were signaling that the main focus of research concerning the Commission was statistical thermodynamics. The animated discussions following most presentations at the Brussels symposium showed that many of the employed assumptions and approaches were matter of contention. This was particularly evident in the opening scientific talk,“The statistical basis of thermodynamics," which explored the foundation of thermodynamic laws on a unique statistical principle [34]. [^78] Guggenheim was also present at the Florence conference, where he gave a talk discussing the Boltzmann factor [33]. Given by the president of the Commission, Edward Guggenheim, the opening talk of the symposium suggested a collective lean towards recognizing statistical mechanics as a vital area for developing an international research community. While in the Brussels’ proceedings the lion’s share is taken up by the thermodynamics of open systems promoted by Prigogine and the locals, in Florence several contributions focused on more fundamental problems of statistical mechanics [13, 43, 33, 8, 41]; this passage accompanied a more basic Commission’s shift in the scientific direction, prefigured by its renaming. As discussed in the Section 4, moving from Commission on Thermodynamic Magnitudes and Notations to Commission on Thermodynamics and Statistical Mechanics had already marked the progression toward assuming a scientific role rather than a purely technical one as previously envisaged. This strategic shift to statistical mechanics, seen as closer to the heart of physics and offering a richer vein of research possibilities, was evident in the content featured at the Florence conference in contrast to the Brussels symposium.
@@
The Commission’s shift toward statistical mechanics was operationalized by the Florence conference through the delineation of the five above-mentioned sub-topics. This strategic focus marked the Commission’s evolution beyond its traditional thermodynamics realm, aiming instead to anchor itself within the physics community by spearheading the exploration of specific, cutting-edge research areas. The commitment to these sub-topics signified an intention not just to serve as a community hub but to actively shape the research landscape. This ambition was realized as the proceedings of the conference showcased: all five sub-topics were comprehensively addressed by the speakers, albeit with a variable number of papers dedicated to each sub-topic. [^79] For the mathematical methods of the configuration integral, see [56]; cooperative phenomena were addressed in [33, 71]; distribution functions were the topic of [54, 67]; the statistical mechanics of irreversible phenomena was explicitly tackled by [18, 41, 82]; and the quantum foundations of statistical mechanics was the topic of Born’s lecture [13]. Such thorough coverage reaffirms the conference’s success in redirecting the attention of the community around the Commission towards statistical mechanics. The opening speech of Kramers (1949) openly reveals the intention to position statistical mechanics as a field with historical roots in the kinetic theory of gases, now poised to incorporate the recent advances in quantum physics and to confront fundamental and outstanding problems, including the derivation of the thermodynamic properties of a system from the interaction among its elementary constituents, starting from the comprehension of ferromagnetism and the condensation of a gas.
@@
Some presentations at the conference have been recognized, a posteriori, as influential on the later developments of what was then considered an emerging field. This is particularly manifest in Lars Onsager’s lecture, titled Statistical Hydrodynamics. His paper published in Supplemento al Nuovo Cimento [62] has amassed over 2,000 citations to date, dwarfing the citation counts of other papers from the proceedings by two orders of magnitude. [^80] This analysis was made using Google Scholar. A research on Scopus confirms the later impact of the paper with a count of 1362 citations from 1970 to 2024 with an evident increase of citations starting from the late 1990s. Onsager’s paper studies the statistics of a system of vortices, that represents the first and one of the most important examples displaying (absolute) negative temperatures. In it, Onsager, without providing a full demonstration, introduced the hypothesis of quantized vortical lines in superfluid helium as a hydrodynamic invariant. Remarkably, at the time, the significance of this work was not fully recognized, just as it is notable that the discovery for which Onsager was awarded the Nobel Prize for Chemistry in 1968 [his reciprocal relations identified in 59, 60] was the focus of Casimir’s presentation at the Florence Conference [18], rather than his own. Only after this conjecture was reformulated by Feynman in the 1950s and, consequently, put to test, the powerful anticipation by Onsager at the Florence conference was fully recognized [27]. [^81] It is also worth noting a short comment of Onsager (him again) to the paper by Rushbrooke (1949, p. 261), where Onsager gives the analytical expression of the magnetization of the two dimensional Ising model, a result which will be published (together with its independent derivation) three years later by C.N. Yang (1952).
@@ 7 The solidification of a tradition from topical conferences to StatPhys
1 Florence (Italy) [^1] International conference on statistical mechanics 1949 16 Boston (USA) 1986
@@ 15 Edinburgh (UK) 1983
8383footnotetext: This table was compiled using data sourced from the IUPAP reports, Series B2aa “General Reports,” Vols.1-3, IuG. The denomination StatPhys was officially introduced in 1977 to designate the Haifa conference as StatPhys13. By assigning this sequence number, the Commission implicitly applied the StatPhys label to all prior meetings in the series, thereby retrospectively recognizing them as part of the StatPhys conference lineage, as shown in the table.
@@
The 1970s witnessed significant transformations concerning the policies and structures of IUPAP commissions and their sponsored conferences. Starting from the 1969 General Assembly in Dubrovnik, IUPAP embarked on a comprehensive revision of policies related to the composition of commissions and the criteria for conference sponsorship. In response to the increasing volume of conference proposals from IUPAP commissions, the Executive Committee of IUPAP introduced a set of evaluation criteria focused on scientific merit, international representation, and organizational solidity. A member of the U.S. national committee, William Havens Jr., further contributed to these deliberations by proposing a taxonomy for classifying conferences into three distinct categories, which received tentative approval during the 13th General Assembly in Dubrovnik. The type-A conferences, called “general conferences," were designed “to provide an overview of the entire field of interest to a Commission, and would normally occur at three-year intervals" with a foreseeable attendance of about 750-1500 scientists. Type-B conferences were called “topical conferences" and were designed to “concentrate on broad sub-fields" within the general area of interest of the Commission. The foreseen attendance was about 300-600 physicists. The smaller Type C conference were called “special conferences" and “would concentrate on much more restricted specialized topics" than the type-B conferences. The expected attendance was in the range of 50-200 physicists. [^84] IUPAP, Report 1970, esp. p. 25, p. 29, pp. 81-83, Series B2aa, Vol. 2, IuG. At the subsequent IUPAP General Assembly, held in Washington D.C. in 1972, new procedural guidelines were established for the composition of commissions and the election of their members. Incidentally, a new identification system for commissions was adopted, denoted as “C. followed by a sequential number reflecting their chronological establishment, a system that remains in use today.” [^85] IUPAP, Report 1973, pp. 14-18, Series B2aa, Vol. 2, IuG. [^86] The full point after C was later eliminated and currently IUPAP commissions are simply denoted by C plus a sequential number, with C3 being the denomination of the Commission on Statistical Physics.
@@
The reorganization within IUPAP had significant implications for the operations of all its commissions. In 1972, the designation of the Commission on Thermodynamics and Statistical Mechanics as C.3 cemented its reputation as the inaugural topical commission of IUPAP, following C.1, the Commission on Finances, and C.2, the general-purpose SUN Commission. Prompted by these general reconfigurations within IUPAP, in 1973 the Commission resolved to inaugurate an award recognizing outstanding achievements in thermodynamics and statistical mechanics. The newly appointed chairman, Canadian physicist Donald D. Betts, announced to the Commission members that the concept of the award was met with enthusiastic endorsement, both within the Commission and among the wider physics community. Betts advocated for the award to take the form of a gold medal rather than a cash prize, arguing that it would not only be economically feasible for the Commission but also offer a lasting symbol of achievement. Following a Commission’s vote on names of founding fathers of the sub-discipline, the award was eventually called the Boltzmann Medal. The inaugural Boltzmann Medal was presented at the 1975 conference on Statistical Mechanics to Kenneth Wilson, recognizing his pioneering work on the renormalization group. This decision came despite Betts’ earlier communication in 1973, which had suggested Lars Onsager as the “obvious choice" [^87] Betts to Members of the Commission on Thermodynamics and Statistical Mechanics, March 9, 1973, see also Betts to Larkin Kerwin, September 19, 1973, Series E3 “Correspondence of the Commissions,” Vol. 2, Folder C 3, IuG. for his contributions to the study of irreversible processes, for which he had already been awarded the Nobel Prize in Chemistry in 1968.888888Lars Onsager died in 1976. It was therefore not possible to take him into consideration as a recipient of the next Boltzmann Medals. Later, in 1982, Kenneth Wilson also won the Nobel Prize in Physics. The same occurred to Giorgio Parisi who received the Boltzmann medal in 1992 and the Nobel prize in 2021.
@@
The second significant shift in the mid-1970s aimed to establish a more systematic scheduling of the Commission on Thermodynamics and Statistical Mechanics’ international conferences, proposing a biennial interval for these conferences starting in 1975.898989Betts to Kerwin, September 19, 1973, Series E3 “Correspondence of the Commissions,” Vol. 2, Folder C 3, IuG. This initiative was in line with the above-mentioned broader initiative to categorize IUPAP-sponsored international conferences, as endorsed during the 1969 IUPAP General Assembly. The conferences organized by Commission C.3 naturally fell into the Type-A category, embodying the essence of a general conference. The decision in the early 1960s to forgo specific titles in favor of a more inclusive approach was exactly in this direction. However, aligning the Commission’s ambition for biennial conferences with IUPAP’s organizational framework proved challenging. The financial and administrative constraints of adhering to a two-year cycle were compounded by IUPAP’s triennial budgetary cycle, which spanned from one General Assembly to the next. As Larkin Kerwin, the IUPAP secretary general, explained to Betts, supporting Type-A conferences biennially within the confines of a three-year budget was problematic. [^90] Kerwin to Betts, October 22, 1973, Series E3, Vol. 2, Folder C 3, IuG. Additionally, during this period, IUPAP officials urged all commissions to reassess their roles in light of recent disciplinary advancements. Commission C.3 was tasked to provide a report in 1977 “examin[ing] the boundaries of its field and to consider whether these should be modified, whether the Commission should be divided, fused with another commission, or dissolved" as well as to reassess its field through the lens of its international conferences. [^91] Kerwin to Herbert B. Callen, October 10, 1975, Series E3, Vol. 2, Folder C 3, IuG. This directive underscored the need for commissions to remain adaptive and responsive to the evolving landscape of their respective disciplines.
@@
This request initiated a self-reflection among the members of Commission C.3 regarding its activities, roles, and functions. In 1977, the chair of the Commission in 1975-78, U.S. statistical physicist Herbert B. Callen, submitted a preliminary report to Kerwin, articulating the Commission members’ perception of both the Commission and the field it served during that time. It is beneficial to quote extensive passages from this report, as they directly illustrate the views of Commission’s members about the Commission itself and its scientific domain within the broader landscape of physics.